40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 19 23:50:44 2013, Jenne, Update, LSC, PRCL locking again 
    Reply  Thu Jun 20 01:13:18 2013, Jenne, Update, LSC, PRCL locking again - no ASC success 
       Reply  Wed Jun 26 04:38:02 2013, Jenne, Update, LSC, PRCL locking again - ASC success PRMI_ASC_yawOnly_powerSpectra_25June2013.pdfPRMI_ASC_yawOnly_25June2013_mag.pdfPRMI_ASC_yawOnly_25June2013_phase.pdfPRMI_ASC_yawOnly_25June2013_coherence.pdfSensMatMeas_26June2013.png
          Reply  Thu Jun 27 17:09:41 2013, Jenne, Update, LSC, PRCL locking again - POP PIT work MeasuredVsModeledASCyaw.png
             Reply  Fri Jun 28 02:12:41 2013, Jenne, Update, LSC, PRCL locking again - ASC work PRMI_ASC_powerSpectra_27June2013.pdfScreenshot-Untitled_Window.png
Message ID: 8753     Entry time: Wed Jun 26 04:38:02 2013     In reply to: 8731     Reply to this: 8767
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRCL locking again - ASC success 

With Rana's help/supervision/suggestions, I have closed the loop on the PRMI ASC servo with the new QPD.  I think I've had it locked for ~30+ minutes now.  It was locked for ~45 minutes, but then the MC momentarily lost lock.  I immediately recovered the PRMI+ASC (after small PRM yaw tweaking, since the ASC isn't triggered yet, so the MC lockloss caused a big yaw step function to go to the PRM, which displayed a bit of hysteresis.).

My biggest problem was that I didn't really understand Koji's servo filter choices, so I wasn't using the right ones / doing good things.  In particular, I need to compensate for the oplev servo filters.  The oplev servo shape is something like ^, so the 1/(1+G) shape is something like =v= (ignoring the lower horizontal lines there).  For tonight, we just turned off the PRM oplevs, but clearly this isn't a permanent solution.  (Although, after Rana went in and roughly centered the PRM oplev, we noticed that turning the oplev on and off doesn't make a huge difference for the PRM....we should investigate why not.  Also, we turned off the FM2 3.2Hz resonant gains in the PRM oplevs, since the Q of those filters is too high, much higher than our actual stacks). 

Rana and I also locked the PRM-ITMY half cavity, and used that beam to realign the beam onto the POP QPD, POP110 PD, and the camera. 

The POP QPD pitch and yaw signals with the half cavity have some noise, that looks like 60Hz crap.  Since this goes away (rather, is much less noticeable) with the regular sideband-locked PRMI, we suspect this is a problem with perhaps the normalization, with the sum very low, and having some noise on it.

Once we had our ASC filters set up (not the 10Hz boost yet though, I think), if I increased the gain from -0.02 to -0.03, we start to get some gain peaking.  With a gain of -0.04, the peak is very noticeable around 250Hz.  We aren't sure where this is coming from, since it shouldn't be coming from the ASC loop.  The UGF of that loop is much lower (I measured it, to check, and the UGF is ~5Hz). Anyhow, this is still a mystery, although the gain of -0.02 holds the cavity pretty well.

I measured the power spectra of the POP QPD pit, yaw, sum, as well as POPDC and POP110I, with the ASC loop on and off (dashed lines are with the loop on.  You can see that the yaw motion as seen on the QPD was reduced by almost 2 orders of magnitude below 1Hz.  It also looks like we can win some more by turning on the equivalent pitch ASC servo (this is also something we see when looking at the dataviewer traces).

I also tried to measure the PRMI sensing matrix, but I get some weird results, even after I double the drive actuation.  I need to be checking whether or not my drive is actually coherent with the error signals that I'm seeing, because right now I'm not sure that I believe things. I'm going to leave that on the to-do list for tomorrow night though.

Next up:

* Engage POP QPD -> pitch loop, copying yaw loop.

* enable ASC triggering

* model PRMI sensing matrix and error signals, bringing one arm into resonance

* Lock the PRMI, and bring the Xarm into IR resonance using the ALS system.

-------------------------------------------------------------------------------------

Here are some numbers and plots from the night:

Right now, I'm locking the LSC with:

MICH LSC with AS55Q, FMs 4 and 5 on, FM 3 is triggered, gain = -40.0, normalized by sqrt(POP110I)*0.1

PRCL LSC with REFL33I, FMs 4 and 5 on, FM 9 is triggered, gain = +2.5, normalized by sqrt(POP110I)*10

(FM3 of MICH and FM9 of PRCL are the same, just in different spots).

The ASC (only POP yaw -> PRM yaw right now) has:

FMs 1,2,5,6 on (1 = integrator [0:0.1], 2 = 3.2 res gain, 5 = [1000,1000:1 and gain of 0.01], 6 = 10Hz boost).  Gain = -0.020,  Limit=5000.

Turn off the input, turn on the output and the gain, clear the histories (to clear out the integrator in FM1), then turn on the input.

PRM oplev is OFF. (need to put in a filter to compensate for it in the ASC servo, but for tonight, we just turned it off.)

We measured the spectra of the POP QPD signals with the ASC loop on and off:

PRMI_ASC_yawOnly_powerSpectra_25June2013.pdf

I also measured the ASC loop (with the PRM oplev still off):

 PRMI_ASC_yawOnly_25June2013_mag.pdf

PRMI_ASC_yawOnly_25June2013_phase.pdf

PRMI_ASC_yawOnly_25June2013_coherence.pdf

(sorry about the separate plots - I can't make DTT give me more than 2 plots on a page at a time right now, so I'm giving up, and just making 3 separate pages)

Weird sensing matrix, unsure if I'm really getting good coherence:

SensMatMeas_26June2013.png

ELOG V3.1.3-